### v1

@@ 2.5 Normalizing Flows (Flow-based Generative Models)
For a multivariate vector x, the objective of normalizing flow is to learn the probability density function $p(\textbf{x})$. The density $p(\textbf{x})$ is transformed into a simple distribution $p(\mathbf{z}_{0})$ (independent multivariate Gaussian distribution). This transformation $f$ is composed of $K$ invertible (bijective) functions $f_{i},i\in\{1,...K\}$. This transformation can be denoted as follows:
@@
what is the meaning of $\circ$
Figure 3: Schematic overview of normalizing flow
@@
$\begin{split}\mathcal{L}(\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\log p(\mathbf{x})\end{split}.$ (2.21)
@@
$\begin{split}&\mathbf{y}_{1:d}=\mathbf{x}_{1:d}\\ &\mathbf{y}_{d+1:D}=\mathbf{x}_{d+1:D}\odot\exp{\left(s(\mathbf{x}_{1:d})\right)}+t(\mathbf{x}_{1:d}),\\ \end{split}$ (2.22)
@@
$\begin{split}&\mathbf{x}_{1:d}=\mathbf{y}_{1:d}\\ &\mathbf{x}_{d+1:D}=\left(\mathbf{y}_{d+1:D}-t\left(\mathbf{y}_{1:d}\right)\right)\odot\exp{\left(-s\left(\mathbf{y}_{1:d}\right)\right)}.\\ \end{split}$ (2.23)
@@
$\begin{split}\frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}=\begin{bmatrix}\mathbbm{I}_{d}&0\\ \frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}&\operatorname{diag}\left(\exp{\left[s(\mathbf{x}_{1:d})\right]}\right)\end{bmatrix},\end{split}$ (2.24)
@@
$\begin{split}\left|\det\left(\frac{\partial y}{\partial x^{\top}}\right)\right|&=\prod_{i=1}^{D}\left[\frac{\partial y}{\partial x^{\top}}\right]_{i,i}=\prod_{i=1}^{d}\exp{\left[s(\mathbf{x}_{1:d})\right]_{i,i}}\\ &=\exp{\left(\sum_{i=1}^{d}\left[s(\mathbf{x}_{1:d})\right]_{i,i}\right)},\\ \end{split}$ (2.25)
@@
$\begin{split}\log\left|\det\left(\frac{\partial y}{\partial x^{\top}}\right)\right|&=\sum_{i=1}^{d}\left[s(\mathbf{x}_{1:d})\right]_{i,i},\\ \end{split}$ (2.26)
@@ 2.6 Score-Based Generative Models
Score-based generative models are a class of generative models that leverage the concept of “score” of a probability density function to generate new data samples. The term score is used in a statistical context, referring to the gradient (or intuitively, slope) of the log-likelihood function with respect to the data, $\mathbf{x}$. Formally, the score is denoted as $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Intuitively, the score tells us how to change $\mathbf{x}$ to increase the probability of $\mathbf{x}$ under the model’s current estimate of the probability density function. A higher score indicates a direction in which the data point $\mathbf{x}$ could be adjusted to make it more likely under the model.
Figure 4: Schematic overview of score-based generative model
@@
Once the score network is trained, data generation can be accomplished through an iterative method known as Langevin dynamics, or Langevin Monte Carlo (Parisi, 1981; Grenander and Miller, 1994). Langevin dynamics is originally formulated to describe the behavior of molecular systems, such as Brownian motion. A more detailed derivation of equations for Langevin dynamics in the generative model setting is presented in Appendix B. Given a fixed step size $\epsilon>0$, and an initial random noise $\mathbf{x}$, the following iterative procedure can be used:
@@ 2.7 Diffusion Models
Figure 5: Schematic overview of diffusion model
@@
Instead of directly training the model with negative log-likelihood, we can consider $q$ as the approximate posterior and use the variational bound on negative log-likelihood to train the model. A detailed derivation is presented in Appendix A and Ho et al. (2020). As a result, we can find an upper bound of the optimization
@@ 3.1.1   Scenario and Synthetic Data Generation
Scenario and synthetic data generation involve using DGMs to produce a wide range of synthetic data. This process includes creating diverse driving scenarios to test and evaluate autonomous vehicles, synthesizing demographic and travel behavior data to represent different populations in studies, and generating synthetic traffic data for training other machine learning models for traffic operation and management. Additionally, DGMs can generate rare or unexpected traffic situations to improve anomaly detection and response strategies. These capabilities allow researchers to simulate various conditions and events which can help enhance the robustness and accuracy of analysis. This section highlights how the DGMs enable comprehensive testing, analysis, and improvement of transportation networks through effective scenario and synthetic data generation.
@@ 3.1.1.1   Driving Scenario Generation
Many studies in this section have focused on driving scenario generation, which involves creating diverse and realistic driving situations to test and evaluate autonomous vehicles and traffic management systems (Ghosh et al., 2016; Yun et al., 2019; Singh et al., 2023; Ye et al., 2022; Jin et al., 2023; Dong et al., 2023; Xu et al., 2023b; Huang et al., 2024; Li et al., 2024d). Most of these studies utilize additional information as conditions for their generative models, such as car-following theory, safety critics, or physical restrictions, to generate more realistic and high-performing models. For example, Chen et al. (2022a) developed a two-step hybrid driving model that combines model-based controllers with Generative Adversarial Imitation Learning (GAIL) for traffic simulation. The model generates high-level driver traits used as parameters for low-level model-based controllers to simulate human-like driving in multi-agent traffic scenarios. Zhong et al. (2023) presented a framework using a conditional diffusion model to generate realistic and controllable traffic simulations. The key benefit of the model is that it allows users to specify trajectory properties, such as reaching a goal or following speed limits while ensuring that the generated trajectories are physically feasible and realistic through the guidance of signal temporal logic rules. Similarly, Sun et al. (2023) introduced a data-driven method DriverSceneGen for generating diverse and realistic driving scenarios to address data limitations. DriverSceneGen employed a diffusion model to create initial driving scenes in a rasterized Bird-Eye-View (BEV) format and used a simulation network to predict multiple future scenarios based on the initial setup. Furthermore, Xu et al. (2023a) developed DiffScene, a diffusion-based framework designed to generate safety-critical driving scenarios for evaluating and enhancing autonomous vehicle safety. DiffScene used a diffusion model to approximate the distribution of low-density areas in traffic data, creating realistic safety-critical situations. These scenarios were further optimized using guided adversarial objectives to maintain both realism and safety-criticality. Additionally, Zhang et al. (2020a) employed the conditional GAIL (cGAIL) framework to study taxi drivers’ decision-making strategies. The method transferred knowledge across different drivers and locations to improve policy accuracy, improving taxi drivers’ income and service quality. These studies have demonstrated the capability of using DGMs in driving scenario generation by providing more diverse and realistic scenarios for traffic management and Autonomous Vehicle development.
3.1.1.2   Population Synthesis
Population synthesis is a critical component of transportation research that involves generating synthetic demographic and travel behavior data to accurately represent populations. This synthetic data is essential for effectively modeling and analyzing transportation systems. Many studies have utilized VAEs or GAN-based models to solve the problems (Borysov et al., 2018; Badu-Marfo et al., 2020; Mensah et al., 2022; Johnsen et al., 2022; Badu-Marfo et al., 2022). For instance, Borysov et al. (2019) presented a VAE-based approach for population synthesis to address common challenges, such as scalability and the issue of sampling zeros in previous methods. The proposed method fits the full joint distribution of high-dimensional data, enabling the generation of more comprehensive synthetic datasets. Validation with a large-scale Danish trip diary dataset demonstrated the method’s effectiveness in enhancing the realism and diversity of synthetic populations. Kim and Bansal (2023) also tackled the challenges of structural zeros and the generation of feasible, diverse synthetic populations. They employed customized loss functions for the VAE and Wasserstein GAN (WGANs) and validated their approach using a large household travel survey dataset from South Korea. Similarly, Garrido et al. (2020) proposed a VAE and WGAN-based method to predict rare feature combinations in population synthesis. These models generate synthetic populations that include rare and diverse feature combinations throughout capturing the joint distribution of high-dimensional attributes. Validation with the Danish National Travel Survey data demonstrated notable improvements in generating realistic and varied synthetic populations.These studies collectively highlight the versatility and robustness of DGMs in population synthesis. They underscore the importance of DGMs in creating realistic and diverse synthetic data, which is crucial for accurate transportation modeling and analysis.

### v3

@@ 2.5 Normalizing Flows (Flow-based Generative Models)
As shown in Figure 3, for a multivariate vector x, the objective of normalizing flow is to learn the probability density function $p(\textbf{x})$. The density $p(\textbf{x})$ is transformed into a simple distribution $p(\mathbf{z}_{0})$ (independent multivariate Gaussian distribution). This transformation $f$ is composed of $K$ invertible (bijective) functions $f_{i},i\in\{1,...K\}$. This transformation can be denoted as follows:
@@
Figure 3: Schematic overview of normalizing flow. Here, $\mathbf{z}_{0}$ is a simple, known distribution (such as a standard Gaussian), $\mathbf{z}_{i}$ represents an intermediate distribution, and $\mathbf{z}_{K}$ denotes the target distribution.
@@
$\mathcal{L}(\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\log p(\mathbf{x}).$ (2.14)
@@
$\begin{split}&\mathbf{y}_{1:d}=\mathbf{x}_{1:d},\\ &\mathbf{y}_{d+1:D}=\mathbf{x}_{d+1:D}\odot\exp{\left(s(\mathbf{x}_{1:d})\right)}+t(\mathbf{x}_{1:d}),\\ \end{split}$ (2.15)
@@
$\begin{split}&\mathbf{x}_{1:d}=\mathbf{y}_{1:d},\\ &\mathbf{x}_{d+1:D}=\left(\mathbf{y}_{d+1:D}-t\left(\mathbf{y}_{1:d}\right)\right)\odot\exp{\left(-s\left(\mathbf{y}_{1:d}\right)\right)}.\\ \end{split}$ (2.16)
@@
$\begin{split}\frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}=\begin{bmatrix}\mathbbm{I}_{d}&0\\ \frac{\partial\mathbf{y}_{d+1:D}}{\partial\mathbf{x}_{1:d}^{\top}}&\operatorname{diag}\left(\exp{\left[s(\mathbf{x}_{1:d})\right]}\right)\end{bmatrix},\end{split}$ (2.17)
@@
$\begin{split}\left|\det\left(\frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}\right)\right|&=\prod_{i=1}^{D}\left[\frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}\right]_{i,i}=\prod_{i=1}^{d}\exp{\left[s(\mathbf{x}_{1:d})\right]_{i,i}}=\exp{\left(\sum_{i=1}^{d}\left[s(\mathbf{x}_{1:d})\right]_{i,i}\right)},\\ \end{split}$ (2.18)
@@
$\begin{split}\log\left|\det\left(\frac{\partial\mathbf{y}}{\partial\mathbf{x}^{\top}}\right)\right|&=\sum_{i=1}^{d}\left[s(\mathbf{x}_{1:d})\right]_{i,i},\\ \end{split}$ (2.19)
@@ 2.6 Score-Based Generative Models
Score-based generative models are a class of generative models that leverage the concept of “score” of a probability density function to generate new data samples. The term score is used in a statistical context, referring to the gradient (or intuitively, slope) of the log-likelihood function with respect to the data, $\mathbf{x}$. Formally, the score is denoted as $\nabla_{\mathbf{x}}\log p(\mathbf{x})$, as shown in Figure 4. Intuitively, the score tells us how to change $\mathbf{x}$ to increase the probability of $\mathbf{x}$ under the model’s current estimate of the probability density function. A higher score indicates a direction in which the data point $\mathbf{x}$ could be adjusted to make it more likely under the model.
Figure 4: Schematic overview of score-based generative model. Here, $\mathbf{x}$ denotes a data sample from the underlying distribution, and $\nabla_{\mathbf{x}}\log p(\mathbf{x})$ represents the score function.
@@
Once the score network is trained, data generation can be accomplished through an iterative method known as Langevin dynamics, or Langevin Monte Carlo (Parisi, 1981; Grenander and Miller, 1994). Langevin dynamics is originally formulated to describe the behavior of molecular systems, such as Brownian motion. A more detailed derivation of equations for Langevin dynamics in the generative model setting is presented in Appendix E. Given a fixed step size $\epsilon>0$, and an initial random noise $\mathbf{x}$, the following iterative procedure can be used:
@@ 2.7 Diffusion Models
Figure 5: Schematic overview of diffusion model. Here, $\mathbf{x}_{0}$ represents the original data, $\mathbf{x}_{t}$ represents the intermediate noisy states, and $\mathbf{x}_{T}$ represents the noise data (e.g., Gaussian distribution).
@@
Instead of directly training the model with negative log-likelihood, we can consider $q$ as the approximate posterior and use the variational bound on negative log-likelihood to train the model. A detailed derivation is presented in Appendix D and Ho et al. (2020). As a result, we can find an upper bound of the optimization
@@
Since the diffusion process follows Gaussian distribution, the true reverse process, $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$, can be assumed to follow a Gaussian distribution if $T$ is sufficiently large, or $T\rightarrow\infty$. Let $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)=\mathcal{N}(\mathbf{x}_{t-1};\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0}),\tilde{\beta}_{t}\mathbbm{I})$. To derive explicit form of $\tilde{\boldsymbol{\mu}}_{t}(\mathbf{x}_{t},\mathbf{x}_{0})$ and $\tilde{\beta}_{t}$, first, we should derive a closed-form equation for sampling $\mathbf{x}_{t}$ at an arbitrary timestep $t$ from Equation (2.25) as shown in Appendix D.
@@ 3.1.1   Scenario and Synthetic Data Generation
Scenario and synthetic data generation involve using DGMs to produce a wide range of synthetic data. This process includes creating diverse driving scenarios to test and evaluate autonomous vehicles, synthesizing demographic and travel behavior data to represent different populations in studies, and generating synthetic traffic data for training other machine learning models for traffic operation and management. Additionally, DGMs can generate rare or unexpected traffic situations to improve anomaly detection and response strategies. These capabilities allow researchers to simulate various conditions and events, which can help enhance the robustness and accuracy of analysis. This section highlights how the DGMs enable comprehensive testing, analysis, and improvement of transportation networks through effective scenario and synthetic data generation. We also categorize recent studies by generation type, task, model used, dataset, and evaluation metrics, as summarized in Table 2.
Table 2: A summary of recent studies in scenario and synthetic data generation using DGM
Generation Type Author Task Model Dataset Evaluation Metrics
Driving Scenario Generation Zhong et al. (2023) Driving Scenarios Diffusion nuScenes Rule-specific Metrics, WD
Sun et al. (2023) Driving Scenarios DriveSceneGen (Diffusion-based) Waymo Motion Data FID, Average Feature Distance, Precision, Recall, F1-Score
Xu et al. (2023a) Driving Scenarios DiffScene (Diffusion-based) CARLA Simulation Rule-specific Metrics, SSPD, FID, DTW, WD, KLD
Niedoba et al. (2024) Driving Scenarios DJINN (Diffusion-based) Argoverse, INTERACTION minADEk, minFDEk, MFDk, minSceneADEk, minSceneFDEk
Population and Activity Synthesis Borysov et al. (2019) Population Synthesis VAE Danish National Travel Survey SRMSE, $R^{2}$, Diversity Measures, Distribution Comparison, Pearson’s Correlation Coefficient
Garrido et al. (2020) Population Synthesis VAE, WGAN Danish National Travel Survey SRMSE, $R^{2}$, Sampling Zeros, Structural Zeros, Pearson’s Correlation Coefficient
Kim and Bansal (2023) Population Synthesis VAE, WGAN Korea Household Travel Survey SRMSE, Recall, Precision, F1-Score, Number of Unique Combinations
Jutras-Dubé et al. (2024) Population Synthesis CTGAN, TVAE American Community Survey SRMSE, Recall, Precision, F1-Score, Sampling Zeros, Structural Zeros
Jeong et al. (2021) Activity-based Modeling VAE Seoul Household Travel Survey Matching Rate, RMSE, Precision, Recall, F1-Score, Accscore
Badu-Marfo et al. (2022) Activity-based Modeling CTGAN 2013 Montreal OD Travel Survey SRMSE, $R^{2}$, Pearson’s Correlation Coefficient, Distribution Comparison
Lee et al. (2025) Activity-based Modeling CollaGAN Seoul Household Travel Survey, Smart Card Data SRMSE, Feasibility, Heterogeneity, F1-Score, Distribution Comparison
Synthetic Data and Model Training Zhou et al. (2023) Urban Flow Generation KSTDiff Urban Flow Data from 4 Cities (Washington D.C., NYC, Beijing, Baltimore) MAE, RMSE, SMAPE, MMD, NRMSE, JSD, Pearson’s Correlation Coefficient
Rong et al. (2023a) OD Generation DiffODGen OD Data from 3 US City (Cook County, New York City, Seattle) RMSE, NRMSE, CPC, JSD, Rate of Nonzero Flows, Accuracy, 1-Recall, Precision
Rong et al. (2023b) OD Generation ODGN (cGAN) OD Data from 8 US City (NYC, LA, Chicago, Houston, SF, Seattle, Washington D.C., Memphis) JSD, RMSE, CPC
Ozturk et al. (2020) RL Agent Training GAN NGSIM Crash Counts, Reward Comparison
Dewi et al. (2022) Recognition Models Training DCGAN Taiwan Prohibitory Signs Dataset MSE, Structural Similarity Index, Intersection over Union, mAP, Detection time, Classification accuracy
Zhu et al. (2022) Prediction Model Training STDGAN PEMSD4, PEMSD7 MAE, RMSE, MAPE
Anomaly Data Generation Cai et al. (2020) Crash Data Augmentation DCGAN Orlando Expressway SR 408 Data AUC, Sensitivity, Specificity
Chen et al. (2022d) Crash Data Augmentation GAM, cGAN, GDAGAN Taiwan Provincial Highway Accident Records Specificity, FPR, Precision, Recall, F1-Score, BA, G-mean
Chen et al. (2024) Crash Data Augmentation CTGAN-RU Washington State Crash Data Specificity, Recall, G-mean, Distribution Consistency, Parameter Recovery
Huo et al. (2021) Extreme Situations Generation T2GAN Sina Weibo Traffic-related Text Data, Beijing Fourth Ring Road Passenger Flow Data MAE, RMSE
@@ 3.1.1.1   Driving Scenario Generation
Many studies in this section have focused on driving scenario generation, which involves creating diverse and realistic driving situations to test and evaluate autonomous vehicles and traffic management systems (Ghosh et al., 2016; Yun et al., 2019; Jin et al., 2023; Xu et al., 2023b; Huang et al., 2024; Li et al., 2024d; Jiang et al., 2024). The use of DGMs is beneficial in this context, as they generate scenarios that overcome the limited variety of real-world data and eliminate the need for costly, time-intensive data collection, thereby enabling scalable testing frameworks. Most of these studies utilize additional information as conditions for their generative models, such as car-following theory, safety critics, or physical restrictions, to generate more realistic and high-performing models. Zhong et al. (2023) presented a framework using a conditional diffusion model to generate realistic and controllable traffic simulations. The key benefit of the model is that it allows users to specify trajectory properties, such as reaching a goal or following speed limits while ensuring that the generated trajectories are physically feasible and realistic through the guidance of signal temporal logic rules. Similarly, Sun et al. (2023) introduced a data-driven method DriverSceneGen for generating diverse and realistic driving scenarios to address data limitations. DriverSceneGen employed a diffusion model to create initial driving scenes in a rasterized Bird-Eye-View (BEV) format and used a simulation network to predict multiple future scenarios based on the initial setup. Furthermore, Xu et al. (2023a) developed DiffScene, a diffusion-based framework designed to generate safety-critical driving scenarios for evaluating and enhancing autonomous vehicle safety. DiffScene used a diffusion model to approximate the distribution of low-density areas in traffic data, creating realistic safety-critical situations. These scenarios were further optimized using guided adversarial objectives to maintain both realism and safety-criticality. Niedoba et al. (2024) introduced DJINN, a diffusion-based generative model that simulates joint traffic scenarios by diffusing the trajectories of all agents in a scene simultaneously. Rather than predicting trajectories for each agent independently, DJINN models the full joint distribution conditioned on flexible observation masks—capturing past, present, or even future state information—and maps context. This approach enables the generation of diverse, realistic traffic scenes, including rare and safety-critical events, while also allowing for test-time guidance and scenario editing. These studies have demonstrated the capability of using DGMs in driving scenario generation by providing more diverse and realistic scenarios for traffic management and Autonomous Vehicle development. Despite these promising developments, evaluating the quality of generated scenarios remains an open problem, as current metrics often fail to fully capture both realism and safety-critical aspects. In particular, there is no appropriate metric to quantitatively compare the distribution of generated driving scenarios to the ground truth. Moreover, existing metrics typically assess only specific features—for instance, verifying whether scenarios obey traffic rules—without capturing the full spectrum of driving realism and safety. Furthermore, current DGM-based frameworks often fail to support long-duration simulations, unlike traditional simulators, which poses an additional challenge. In addition, current studies sometimes fall short in replicating extreme scenarios—such as sudden braking or rapid lane changes—that are critical in highly dynamic, localized environments. This gap underscores the need for future research to develop comprehensive, unified evaluation metrics and more robust simulation platforms that can support extended runs, thereby facilitating rigorous and reproducible testing of various driving systems.
3.1.1.2    Population and Activity Synthesis
Population synthesis is a critical component of transportation research that involves the generation of synthetic demographic and travel behavior data to accurately represent populations. The primary challenge is to capture the full joint distribution of the population while generating new, diverse samples that include rare or unobserved events. Accurately modeling this joint distribution becomes increasingly complex as the number of attributes grows. Simultaneously, during the sampling stage, the exponential increase in possible attribute combinations leads to the sampling zero problem, making it difficult to generate diverse samples that include rare or unobserved events. Recent advances in DGMs, such as VAEs and GANs, offer promising solutions by approximating high-dimensional joint distributions and ensuring sample diversity. Several studies have demonstrated that these approaches can outperform traditional methods such as Gibbs sampling and Bayesian Networks(Borysov et al., 2018; Badu-Marfo et al., 2020; Mensah et al., 2022; Johnsen et al., 2022). For instance, Borysov et al. (2019) employed a VAE to learn latent representations of agents’ attributes, thereby generating synthetic micro-agents that maintain the statistical properties of the original population. Similarly, Garrido et al. (2020) compared VAE and WGAN methods for recovering sampling zeros in high-dimensional travel survey data and found that while WGANs can achieve higher predictive accuracy, VAEs tend to generate more diverse populations. Correspondingly, Kim and Bansal (2023) introduced custom loss functions for both VAE and WGAN to better handle structural-zeros issues, demonstrating improvements in both feasibility and diversity on a large-scale South Korean Household Travel Survey (HTS). Recently, Jutras-Dubé et al. (2024) introduced a copula-based generative framework that generates synthetic data for a target population using only known marginal totals with a sample from a similar population. The proposed method combines copula theory with DGMs (Conditional Tabular GAN (CTGAN) and Tabular VAE (TVAE)) to decouple dependency structures from marginal distributions, ensuring that the synthesized data accurately reflects both the structural relationships and the target population’s marginals. Overall, these advances highlight the versatility of DGMs in addressing the complexities of population synthesis, balancing between achieving predictive accuracy and generating diverse, representative samples.
While synthetic populations form a strong base for transportation analysis, capturing dynamic human mobility requires Activity-Based Modeling (ABM) to simulate sequences of individual activities and predict their interconnections (Liao et al., 2024). The challenge in ABM lies in collecting sufficient data to represent the joint distribution of individual attributes—such as income, job type, gender, and age—which becomes even more complex when incorporating spatial dimensions. Adding spatial attributes increases the likelihood of rare combinations, thereby exacerbating the sampling zero problem. Recent advances in DGMs have been applied to address these challenges (Kim et al., 2022; Jeong et al., 2021; Badu-Marfo et al., 2022; Lee et al., 2025). For example, Jeong et al. (2021) integrated a VAE with a Hidden Markov Model (HMM) to infer and synthesize human activity sequences from mobile data. By embedding high-dimensional, mixed discrete, and continuous features into a lower-dimensional latent space, the proposed VAE-HMM mitigates the challenges of inconsistent clustering and overfitting commonly encountered in traditional HMMs. In another study, Badu-Marfo et al. (2022) proposed a dual-GAN structure, Composite Travel GAN (CTGAN), which simultaneously generates socioeconomic attributes and sequential mobility data learned from the joint distribution of both tabular attributes and sequential trip chain locations data. More recently, Lee et al. (2025) proposed Collaborative GAN (CollaGAN), a GAN-based data fusion method that combines household travel surveys with smart card data. CollaGAN utilizes a semi-supervised variational embedding to harmonize the two datasets within a shared, low-dimensional latent space, employing multiple loss functions—including boundary loss and expert-designed constraints—and dual discriminators to enhance both the feasibility and heterogeneity of the generated activity schedules. These advances demonstrate the potential of DGMs to overcome the challenges of data sparsity and high-dimensional complexity in ABM, ultimately improving the accuracy and diversity of synthesized activity patterns.
However, current DGMs for population and activity synthesis face several challenges. First, these models often suffer from a lack of interpretability, making it difficult to understand the underlying decision processes. Additionally, the characteristics of HTS including a relatively small data size compared to the number of attributes, infrequent updates, and lengthy collection durations, limit the effective utilization of these models. Moreover, the sequential nature of the generated outputs can lead to logically inconsistent travel patterns—such as implausible transitions between transportation modes or mismatches between socioeconomic status and residential characteristics—that necessitate extensive post-processing corrections. Addressing these gaps is essential for advancing DGMs toward more realistic, transparent and policy-relevant transportation models.